# Does Data Contamination Detection Work (Well) for LLMs? A Survey and Evaluation on Detection Assumptions

### v1

@@ A.3 Case Study for Instance Similarity
To assess the applicability of instance similarity-based detection approaches, we analyzed how frequently their requirements are met. We reviewed the top 10 models from the Vellum LLM leaderboard [^8] https://www.vellum.ai/llm-leaderboard#model-comparison. Accessed on Oct 6, 2024.. As demonstrated in Table 5, none of the models fulfilled the most basic, requirement 1. However, some models disclose their cut-off date for collecting training data Achiam et al. (2023); OpenAI .
@@
Table 5: None of the top 10 LMs, in the LLM Leaderboard by Vellum meet the requirements of disclosing pre-training corpora (Requirement 1).
@@ A.4 Entropy Calculation
In casual language modeling, we consider an LM $M$ with a prefix sequence $x_{p}$. We deploy the token entropy to measure the certainty of the next token generation. We first find the next token probability distribution over the vocabulary. Given each Pythia model has more than 50k tokens in the vocabulary Biderman et al. (2023), the entropy is measured over the top $k$ most likely tokens (Entropy $k$). We first normalize the probability of he $k$ tokens:
$P_{i}^{\prime}(M(x_{p}))=\frac{P_{i}(M(x_{p}))}{\sum_{i=1}^{k}P_{i}(M(x_{p}))}$ (10)
The entropy is then calculated over the normalized probability among the $k$ tokens.
$\begin{split}\text{Entropy}_{k}&(M,x_{p})\\ &=-\sum_{i=1}^{k}P_{i}^{\prime}(M(x_{p}))\log P_{i}^{\prime}(M(x_{p}))\end{split}$ (11)
@@ A.5.1 Within-Domain Detection with Different LMs
In this section, we present the detection AUCs for other sizes of Pythia models: Pythia-70m (Table 6), Pythia-160m (Table 7), Pythia-410m (Table 8), Pythia-1.4b (Table 9), Pythia-2.8b (Table 10), Pythia-12b (Table 11). Similar to the results in Table 2, we observed close-to-random performance in the detection AUCs.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 10.8$\pm$25.2 13.1$\pm$5.4 17.2$\pm$9.2 16.0$\pm$7.6 137.4$\pm$1682.7 29.1$\pm$15.6 26.2$\pm$7.5 43.3$\pm$49.5 37.3$\pm$46.5 42.0$\pm$24.1 50.0$\pm$44.6
Unseen 10.5$\pm$19.6 13.7$\pm$13.4 18.6$\pm$34.0 15.8$\pm$7.4 38.9$\pm$44.7 29.2$\pm$22.7 26.1$\pm$8.7 41.7$\pm$31.5 36.9$\pm$37.1 40.8$\pm$21.0 49.8$\pm$36.1
Table 6: Average contamination detection AUC for the Pythia-70m model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 7.1$\pm$15.6 8.5$\pm$3.4 11.8$\pm$5.9 11.1$\pm$5.0 25.1$\pm$34.8 19.0$\pm$10.3 20.0$\pm$7.2 27.3$\pm$23.9 25.2$\pm$29.9 28.5$\pm$15.2 33.1$\pm$29.8
Unseen 7.4$\pm$18.8 8.9$\pm$9.7 12.1$\pm$5.1 11.0$\pm$5.0 25.5$\pm$30.8 18.8$\pm$11.3 19.9$\pm$6.4 26.9$\pm$19.3 24.8$\pm$25.2 27.9$\pm$14.2 32.9$\pm$24.3
Table 7: Average contamination detection AUC for the Pythia-160m model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 4.8$\pm$11.4 5.9$\pm$2.2 8.3$\pm$4.7 7.9$\pm$3.3 16.1$\pm$22.9 13.0$\pm$7.0 14.9$\pm$4.0 17.7$\pm$14.4 17.0$\pm$21.4 19.5$\pm$10.8 22.4$\pm$20.7
Unseen 4.7$\pm$8.7 6.1$\pm$6.0 8.5$\pm$3.6 7.9$\pm$3.4 16.2$\pm$19.5 12.9$\pm$7.3 14.9$\pm$4.5 17.5$\pm$13.4 16.6$\pm$18.4 19.2$\pm$10.3 22.3$\pm$17.3
Table 8: Average contamination detection AUC for the Pythia-410m model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 3.7$\pm$8.0 4.7$\pm$1.7 6.7$\pm$2.9 6.5$\pm$2.5 11.2$\pm$14.9 10.3$\pm$5.4 12.2$\pm$3.3 13.6$\pm$11.5 13.5$\pm$16.5 15.1$\pm$8.1 17.3$\pm$15.3
Unseen 3.7$\pm$6.9 4.9$\pm$4.2 6.8$\pm$2.7 6.5$\pm$2.7 11.4$\pm$14.8 10.2$\pm$5.4 12.2$\pm$3.7 13.5$\pm$11.6 13.0$\pm$14.9 14.8$\pm$7.4 17.2$\pm$13.2
Table 9: Average contamination detection AUC for the Pythia-1.4b model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 3.3$\pm$6.6 4.2$\pm$1.5 6.1$\pm$2.8 6.0$\pm$2.2 9.4$\pm$13.3 9.3$\pm$4.8 10.8$\pm$3.2 11.9$\pm$9.4 12.1$\pm$15.3 13.1$\pm$6.6 15.4$\pm$13.8
Unseen 3.3$\pm$6.5 4.3$\pm$3.5 6.2$\pm$2.5 6.0$\pm$2.4 9.5$\pm$12.4 9.2$\pm$4.8 10.8$\pm$3.5 11.8$\pm$9.0 11.7$\pm$14.2 12.9$\pm$6.0 15.2$\pm$11.4
Table 10: Average contamination detection AUC for the Pythia-2.8b model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens.
Assumptions & Metric Github
Free-
Law
@@
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
@@
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 2.8$\pm$5.4 3.8$\pm$1.3 5.4$\pm$2.6 5.4$\pm$1.9 6.8$\pm$8.9 8.2$\pm$4.0 8.6$\pm$3.5 10.0$\pm$7.9 10.3$\pm$12.6 11.1$\pm$5.3 13.1$\pm$11.1
Unseen 2.9$\pm$6.0 3.8$\pm$2.7 5.4$\pm$2.0 5.4$\pm$2.1 7.0$\pm$9.4 8.1$\pm$3.9 8.6$\pm$3.7 9.9$\pm$8.0 10.1$\pm$12.5 10.8$\pm$4.8 13.0$\pm$9.7

### v3

@@ A.3 Case Study for Instance Similarity
To assess the applicability of instance similarity-based detection approaches (see Section 4.1), we analyzed how frequently their requirements, R1 and R2, are met. We reviewed the top 10 models from the Vellum LLM leaderboard [^6] https://www.vellum.ai/llm-leaderboard#model-comparison. Accessed on Oct 6, 2024.. As demonstrated in Table 6, none of the models fulfilled R1, the most basic requirement. However, some models disclose their cut-off date for collecting training data Achiam et al. (2023); OpenAI .
@@
Table 6: None of the top 10 LMs, in the LLM Leaderboard by Vellum meet the requirements of disclosing pre-training corpora (R1).
@@ A.4 Entropy Calculation
In this section, we explain the procedure for verifying Assumption A6. In the context of casual language modeling, we consider an LM $M$ with a given prefix sequence $x_{p}$. Assumption A6 assumes that given $x_{p}$, if $M$ has seen an instance with the same prefix, it will generate similar responses, regardless of the sampling strategy used. Since the verification of this assumption can be influenced by various sampling strategies, we quantify the variance in the model’s output by measuring the entropy of the token probabilities, which indicates the model’s certainty about the next token generation.
To do this, we first compute the probability distribution of the next token over the model’s vocabulary. Given that LLMs may contain vocabularies with over 50,000 tokens Biderman et al. (2023), most tokens have a very low likelihood of being sampled. Therefore, we focus on the Entropy among the top $k$ most likely tokens (Entropy $k$).
At every token position $x_{p}$, we calculate the entropy based on the probabilities of the top $k$ tokens, using the following formula:
$\begin{split}\text{Entropy}_{k}&(M,x_{p})\\ &=-\sum_{i=1}^{k}P_{i}(M(x_{p}))\log P_{i}(M(x_{p}))\end{split}$ (10)
Given an instance $x$ with $N$ tokens, the Entropy $k$ for $x$ is the average $\text{Entropy}_{k}(M,x_{p})$ across all tokens $x_{p}$ in $x$:
$\text{Entropy}_{k}(M,x)=(\sum_{p=1}^{N}\text{Entropy}_{k}(M,x_{p}))/N$ (11)
@@ A.5.1 Within-Domain Detection with Different LMs
In this section, we present the detailed detection AUCs for all models: (1) different sizes of Pythia Models: Pythia-70m (Table 7), Pythia-160m (Table 8), Pythia-410m (Table 9), Pythia-1.4b (Table 10), Pythia-2.8b (Table 11), Pythia-6.9b (Table 12), Pythia-12b (Table 13); (2) OLMo-2-7B (Table 14); and (3) BioMistral-NLU-7B (Table 15).
Similar to the results in Table 3, we observed close-to-random performance in the detection AUCs for all Pythia models and dataset domains.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.4 49.3 50.5 51.3 51.7 47.4 48.9 52.2 49.9
PPL_100 50.3 50.2 51.5 50.8 50.5 46.2 48.1 52.0 50.8
PPL_200 50.3 50.1 53.5 50.5 49.9 44.3 51.1 50.8 51.5
Min 5% token 51.7 49.9 49.9 50.6 48.9 45.0 50.4 49.9 51.1
Min 15% token 51.5 49.5 51.0 50.5 49.5 45.4 50.3 49.6 51.6
Min 25% token 51.4 50.1 51.0 50.9 49.9 45.5 49.5 49.8 51.3
A4 Mem 5 47.6 49.2 49.5 51.5 50.8 48.8 49.5 50.0 51.2
Mem 15 49.6 49.5 48.7 50.1 50.4 49.0 49.3 50.4 51.2
Mem 25 49.6 49.9 48.3 50.6 50.3 48.2 49.6 49.6 51.0
A6 Entropy 5 50.8 49.7 48.2 52.1 49.9 47.4 49.3 50.4 50.2
Entropy 15 50.6 49.8 48.9 52.4 49.4 47.5 49.3 50.5 50.3
Entropy 25 50.6 49.9 49.2 52.4 49.0 47.7 49.3 50.2 50.2
Average AUC 50.1 49.7 49.8 51.2 50.1 47.3 49.4 50.3 50.9
PPL_200 Seen 11.1$\pm$12.7 13.1$\pm$8.9 29.5$\pm$21.4 28.9$\pm$13.5 46.0$\pm$23.5 36.7$\pm$21.1 42.9$\pm$16.7 45.3$\pm$41.3 50.2$\pm$33.7
Unseen 11.5$\pm$20.9 14.2$\pm$25.3 31.7$\pm$22.4 29.0$\pm$13.1 45.4$\pm$22.5 33.7$\pm$18.2 43.7$\pm$18.8 44.3$\pm$28.7 51.5$\pm$35.6
Table 7: Average contamination detection AUC for the pythia-70m model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.7 49.0 50.3 51.2 51.9 47.6 48.6 52.3 50.0
PPL_100 50.3 49.4 51.2 51.0 50.4 46.1 47.6 51.7 50.7
PPL_200 50.1 49.4 53.0 50.7 49.6 44.2 50.5 50.7 51.7
Min 5% token 51.9 49.9 51.0 50.5 48.6 45.0 48.6 49.4 51.3
Min 15% token 51.4 49.8 50.9 50.8 49.5 45.1 50.1 49.3 51.4
Min 25% token 51.3 49.7 51.3 51.3 49.6 45.6 49.4 49.2 51.4
A4 Mem 5 48.3 49.1 50.4 52.8 51.2 52.6 50.1 49.9 51.1
Mem 15 48.2 49.2 50.1 51.4 51.5 48.0 49.1 49.3 50.0
Mem 25 48.4 49.0 49.9 50.5 51.4 46.2 49.2 49.4 50.2
A6 Entropy 5 51.3 49.1 48.9 52.1 49.8 47.0 48.8 50.1 50.6
Entropy 15 51.1 49.3 49.4 52.2 49.7 47.5 48.9 50.1 50.6
Entropy 25 51.0 49.4 49.6 52.1 49.4 47.7 48.9 50.0 50.6
Average AUC 50.1 49.3 50.4 51.5 50.3 47.7 49.2 49.9 50.8
PPL_200 Seen 7.4$\pm$8.5 8.2$\pm$5.6 18.8$\pm$13.5 18.6$\pm$8.6 29.9$\pm$31.5 45.8$\pm$598.9 29.0$\pm$11.0 30.9$\pm$28.2 33.3$\pm$25.3
Unseen 7.6$\pm$13.5 8.8$\pm$16.7 20.3$\pm$14.8 18.7$\pm$8.3 28.5$\pm$12.8 24.8$\pm$11.4 29.6$\pm$13.1 30.4$\pm$23.0 34.4$\pm$25.9
Table 8: Average contamination detection AUC for the pythia-160m model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.5 49.5 50.1 50.7 51.0 47.4 48.8 51.5 50.0
PPL_100 50.5 49.2 51.3 50.9 49.7 45.5 47.6 51.0 50.3
PPL_200 50.3 49.3 53.0 50.6 48.8 44.6 50.9 50.2 52.0
Min 5% token 52.2 49.8 50.3 50.7 49.0 45.5 49.3 48.9 51.4
Min 15% token 51.6 49.6 50.9 51.0 49.3 46.0 50.2 49.0 51.2
Min 25% token 51.4 49.4 51.0 51.3 48.9 46.4 50.2 48.9 51.0
A4 Mem 5 47.5 49.2 51.3 53.0 50.5 50.3 49.4 49.7 50.3
Mem 15 47.6 49.2 51.3 52.6 51.8 50.3 49.7 50.1 50.4
Mem 25 48.2 49.2 51.1 51.2 51.9 52.6 50.5 49.2 51.2
A6 Entropy 5 51.2 49.1 49.4 52.2 50.5 47.4 49.0 49.5 50.2
Entropy 15 51.0 49.3 49.9 52.0 49.7 48.0 48.9 49.5 50.2
Entropy 25 50.9 49.4 50.1 51.9 49.4 48.2 48.9 49.2 50.3
Average AUC 50.0 49.4 50.8 51.9 50.1 48.4 49.5 49.7 50.7
PPL_200 Seen 4.7$\pm$5.0 5.5$\pm$3.4 11.8$\pm$8.2 12.7$\pm$6.0 18.8$\pm$9.7 20.0$\pm$9.1 19.8$\pm$7.5 20.6$\pm$22.2 22.3$\pm$17.0
Unseen 4.8$\pm$7.2 5.9$\pm$11.4 12.9$\pm$9.6 12.7$\pm$5.7 18.2$\pm$8.3 18.5$\pm$7.9 20.4$\pm$9.3 20.0$\pm$14.8 23.0$\pm$17.8
Table 9: Average contamination detection AUC for the pythia-410m model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.5 49.0 49.8 50.7 50.7 48.1 48.5 51.0 49.8
PPL_100 50.7 49.5 50.9 51.0 49.2 45.6 47.4 50.9 50.2
PPL_200 50.8 49.8 51.9 50.9 48.7 44.4 50.9 49.9 52.0
Min 5% token 51.8 50.5 49.7 51.0 49.2 46.2 48.9 48.4 51.1
Min 15% token 51.5 49.7 50.3 51.3 49.6 47.0 50.1 48.7 51.1
Min 25% token 51.4 49.4 50.5 51.5 49.1 47.8 50.0 48.6 51.2
A4 Mem 5 48.8 49.1 50.9 53.0 51.0 51.5 49.6 49.1 50.5
Mem 15 48.0 49.5 50.8 51.4 51.2 51.7 49.8 48.9 50.4
Mem 25 48.7 49.5 51.1 51.0 51.5 50.9 49.7 48.7 51.0
A6 Entropy 5 51.2 49.0 49.9 52.0 49.9 46.6 48.7 49.3 50.2
Entropy 15 51.0 49.1 50.1 51.9 49.5 48.0 48.7 48.9 50.2
Entropy 25 51.0 49.2 50.1 51.8 49.3 48.5 48.8 48.7 50.3
Average AUC 50.2 49.4 50.5 51.7 50.1 48.6 49.2 49.1 50.6
PPL_200 Seen 3.6$\pm$3.8 4.4$\pm$2.4 8.5$\pm$5.9 9.8$\pm$4.6 14.2$\pm$7.4 16.0$\pm$6.9 15.3$\pm$5.7 16.3$\pm$19.2 17.1$\pm$12.5
Unseen 3.6$\pm$4.9 4.7$\pm$8.6 9.3$\pm$7.3 9.9$\pm$4.4 13.7$\pm$6.4 14.8$\pm$6.1 15.7$\pm$7.2 15.7$\pm$12.3 17.8$\pm$13.9
Table 10: Average contamination detection AUC for the pythia-1.4b model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.4 48.6 49.4 50.7 50.5 47.8 48.0 51.1 49.9
PPL_100 50.5 49.2 50.6 50.9 49.0 45.9 47.5 51.0 50.4
PPL_200 50.6 49.6 51.7 50.8 48.6 45.3 50.8 49.9 52.2
Min 5% token 51.6 49.8 49.9 51.2 49.5 47.2 48.2 48.6 51.6
Min 15% token 51.5 49.6 50.3 51.4 49.5 48.1 49.4 48.7 51.3
Min 25% token 51.2 49.4 50.1 51.4 48.9 48.8 49.7 48.6 51.0
A4 Mem 5 48.7 49.1 50.7 52.8 50.9 51.5 49.5 50.2 50.9
Mem 15 48.2 48.9 50.4 51.0 51.3 50.4 48.9 49.7 50.2
Mem 25 48.9 48.5 50.5 50.8 51.4 50.0 48.7 49.3 50.2
A6 Entropy 5 50.9 49.1 49.5 51.9 50.0 47.5 48.9 49.0 50.6
Entropy 15 50.8 49.2 49.8 51.9 49.5 48.7 49.0 48.9 50.5
Entropy 25 50.8 49.2 50.0 51.8 49.3 49.1 49.0 48.8 50.6
Average AUC 50.1 49.2 50.2 51.6 50.0 48.8 49.0 49.4 50.8
PPL_200 Seen 3.2$\pm$3.3 3.9$\pm$2.1 7.2$\pm$5.1 8.7$\pm$4.1 12.5$\pm$6.5 14.0$\pm$6.2 13.3$\pm$4.9 14.4$\pm$17.3 15.0$\pm$10.2
Unseen 3.2$\pm$4.4 4.2$\pm$7.8 7.9$\pm$6.3 8.7$\pm$3.9 12.1$\pm$5.8 13.1$\pm$5.5 13.6$\pm$6.1 13.9$\pm$11.3 15.7$\pm$12.2
Table 11: Average contamination detection AUC for the pythia-2.8b model, under different domains within the Pile dataset. ‘PPL_200’ represents the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
Assumptions & Metric Github FreeLaw
Enron-
Emails
@@
A1 PPL_50 49.7 48.9 49.5 50.7 50.3 47.7 48.5 50.7 49.8
PPL_100 50.7 49.4 50.3 51.1 48.8 46.2 47.4 50.4 50.4
PPL_200 50.8 49.4 51.1 50.9 48.4 46.7 50.7 49.6 52.1
Min 5% token 51.7 49.8 49.2 51.3 49.7 48.2 47.8 48.5 50.9
Min 15% token 51.6 49.7 49.8 51.3 49.1 49.5 49.7 48.6 51.1
Min 25% token 51.3 49.3 50.0 51.4 48.8 50.2 49.8 48.5 51.1
A4 Mem 5 49.2 48.7 50.6 52.7 50.5 50.0 49.9 49.1 50.8
Mem 15 48.9 48.9 51.0 51.9 51.8 50.6 49.2 48.6 50.7
Mem 25 49.6 48.8 51.1 51.6 51.1 50.3 49.5 48.2 51.3
A6 Entropy 5 51.0 49.1 49.5 52.0 49.7 48.4 49.1 49.3 50.9
Entropy 15 50.8 49.1 49.7 52.0 49.3 49.6 49.0 49.0 50.9
Entropy 25 50.8 49.2 49.8 51.9 49.1 50.0 49.0 48.9 51.0
Average AUC 50.3 49.2 50.2 51.8 49.9 49.4 49.3 49.0 50.9
PPL_200 Seen 2.8$\pm$3.0 3.6$\pm$1.9 6.1$\pm$4.4 7.9$\pm$3.7 11.2$\pm$5.8 12.2$\pm$5.9 12.2$\pm$4.5 13.2$\pm$16.0 13.7$\pm$9.1
Unseen 2.9$\pm$4.1 3.8$\pm$6.2 6.7$\pm$5.6 8.0$\pm$3.6 10.9$\pm$5.3 11.5$\pm$5.1 12.4$\pm$5.4 12.7$\pm$10.6 14.3$\pm$11.0